OpenStudy (anonymous):

explain how you can determine what number goes in the numerator and what number in the denomenator for a percent problem

OpenStudy (skullpatrol):

What does the word "percent" mean?

OpenStudy (anonymous):

amount?

OpenStudy (skullpatrol):

No, not directly...

OpenStudy (skullpatrol):

"percent" is TWO words "per" and "cent" right?

OpenStudy (skullpatrol):

cent means 100, as in 100 cents = 1 dollar

OpenStudy (skullpatrol):

while "per" refers to the denominator of a quotient, right?

OpenStudy (skullpatrol):

Recall: price per item :)

OpenStudy (skullpatrol):

per cent: in or for every hundered

OpenStudy (anonymous):

so for example \(3\%\) means \(3\) out of \(100\) or \(\frac{3}{100}=0.03\)

OpenStudy (anonymous):

and \(20\%\) is \(\frac{20}{100}=\frac{1}{5}=0.2\)

OpenStudy (anonymous):

it is not exactly clear what the question is, but i hope those examples helped